OpenStudy (anonymous):

The average of two numbers is 86. The average of four different numbers is 92. What is the weighted mean of all the numbers together?

OpenStudy (anonymous):

since it doesnt matter what the numbers are, maybe the best way to do it is to pick numbers, 2 numbers are 86 and 4 numbers are 92 then the average would be \[\frac{2\times 86+4\times 92}{6}\]